# Gaussian On-the-Fly Splatting: A Progressive Framework for Robust Near Real-Time 3DGS Optimization

### v1

@@ V-C Ablation Studies
To more explicitly assess the efficacy of each component implemented in our method, we conduct an extensive ablation study as follows. Based on the Palace dataset, we first process it with our full On-the-Fly GS and then sequentially switch off one of the following training strategies: 1) Image Weighting and Training Iteration Allocation (No_IW). 2) Global Optimization (No_GO). 3) Load Balancing (No_Load). 4) Splatting-based Parallelism in Backpropagation (No_SPB). In particular, the cost time in this context refers to the total time required for 3DGS model training after the On-the-Fly SfM completes the pose estimation for each newly fly-in image. The corresponding quantitative results are presented in Table IV and Fig. 6.
TABLE IV: Ablation experiment of the proposal components in the On-the-Fly GS method.
Methods Time Cost (min) Video Memory (MB) PSNR SSIM LPIPS
No_IW 11.21 123.7 21.04 0.83 0.36
No_GO 12.96 271.13 16.62 0.69 0.48
No_Load 14.02 400.47 21.03 0.84 0.33
No_SPB 18.29 182.96 21.21 0.86 0.36
Ours 11.19 173.94 21.24 0.86 0.36
Based on Table IV and Fig. 6, switching off any of the proposed components can negatively impact the 3DGS training, validating their corresponding effectiveness: 1) Image Weighting and Training Iteration Allocation improves the final rendering quality and also slightly reduces time consumption. 2) Global Optimization reduces the memory and time consumption while enhancing model quality. 3) Load Balancing primarily works on memory saving, whereas Splatting-based Parallelism in Backpropagation accelerates training speed. Overall, these ablation study results demonstrate the effectiveness of all the proposed strategies.
Figure 6: Rendering Results of Ablation Studies
@@ VI CONCLUSION
In this paper, we propose On-the-Fly GS, a novel near real-time 3DGS training framework via a progressive optimization manner, which can dynamically output 3DGS field during image capturing. First, we introduce a method for weighting images and allocating different training iterations to each newly captured image during local optimization, enabling fast training for the newly expanded region. Second, we present a new learning rate updating solution based on image rendering quality, adaptively calculating the learning rate for each training iteration based on the training progress of the current image and its neighboring connected images. Finally, inspired by Adr-Gaussian and Taming 3DGS, we further control the number of Gaussians and improve training efficiency. The conducted experiments demonstrate that our On-the-Fly GS successfully achieves near real-time 3DGS training and considerable rendering performance, even when the input images lack spatial and temporal continuity.

### v2

@@ V-C Ablation Studies
TABLE III: Ablation experiment of the proposal components in the On-the-Fly GS method.
Methods Ad-Opt Time (s) Model Size (MB) PSNR SSIM LPIPS
No_FU 110.31 541.1 30.35 0.88 0.21
No_IW 116.76 575.7 31.99 0.91 0.15
No_GO 161.83 860.3 32.17 0.92 0.13
No_Load 252.04 1315.4 34.13 0.95 0.09
No_SPB 213.59 674.5 33.14 0.93 0.11
Ours 105.97 672.3 33.29 0.93 0.11
To explore the efficacy of each component proposed in this work, an extensive ablation study is conducted. Based on the npu dataset, we first turn on all the proposed component (as Ours) and then sequentially switch off one of the following strategies: 1) 3DGS Field Update (No_FU). 2) Image Weighting and Training Iteration Allocation (No_IW). 3) Semi-Global Optimization (No_GO): only train new fly-in image and its neighbors for $N_{I}=200$ iterations, and no training for the very previous images. 4) Load Balancing (No_Load). 5) Splatting-based Parallelism in Backpropagation (No_SPB).
The corresponding quantitative results are presented in Table III. We can see that, removing any of the proposed components can negatively impact the 3DGS training, proving their corresponding effectiveness: 1) Disabling 3DGS Field Update prevents the addition of new Gaussians during scene expansion, resulting in few Gaussians allocated in the newly observed region. 2) Image Weighting and Training Iteration Allocation guide the training process to focus more on the newly expanded region. The optimization of the newly observed region proceeds slowly without this strategy. 3) Removing Global Optimization makes the 3DGS training over-emphasize the optimization on newly observed regions, leading to an excessive accumulation of Gaussians while neglecting rendering quality in previous regions. 4) Load Balancing optimizes the spatial distribution of Gaussians and removes redundant ones, effectively reducing model size while improving training efficiency. 5) Splatting-based Parallelism in Backpropagation accelerates the optimization of the 3DGS field by significantly accelerating the backpropagation operation.
@@ VI CONCLUSION
In this paper, we propose On-the-Fly GS, a novel near real-time 3DGS training framework via a progressive optimization manner, which can dynamically output 3DGS field during image capturing. First, we introduce an self-adaptive method for allocating different training iterations to each newly captured image and its neighboring images during 3DGS optimization, enabling fast training for the newly expanded region. Second, we present a new learning rate updating solution based on image rendering quality, adaptively estimating the learning rate for each training iteration based on the training progress of the current image and its neighboring connected images. Finally, inspired by Adr-Gaussian and Taming 3DGS, we further control the number of Gaussians and improve the training efficiency. The experimental results demonstrate that our On-the-Fly GS can successfully achieve near real-time 3DGS training and considerable rendering performance, even when the input images lack temporal and spatial continuity or are not accompanied by depth. Compared to existing SLAM-based 3DGS methods, On-the-Fly GS exhibits greater versatility and is feasible for a wider range of practical applications.